Ida Sidha Karya Company is a family-owned company located in the village of Gianyar on the island of Bali in Indonesia. The company produces a handcrafted Balinese musical instrument called a gamelan that is similar to a xylophone. The gamelans are sold to a Canadian importer for $877. Selected data for the company’s operations last year follow:

  Units in beginning inventory 0
  Units produced 11,000
  Units sold 7,000
  Units in ending inventory 4,000
  Variable costs per unit:
       Direct materials $ 290
       Direct labour $ 340
       Variable manufacturing overhead $ 60
       Variable selling and administrative $ 15
  Fixed costs:
       Fixed manufacturing overhead $ 850,000
       Fixed selling and administrative $ 910,000

Required:

1. Assume that the company uses absorption costing. Compute the unit product cost for one gamelan.

2. Assume that the company uses variable costing. Compute the unit product cost for one gamelan.
Answer #1

Requirement 1

Unit Product Cost for One Gameplan as per Absorption Costing

Unit Product Cost = Direct Materials Cost per Unit + Direct Labour Cost per Unit + Variable Manufacturing Overhead per Unit + Fixed Manufacturing Overhead per Unit

Direct Materials Cost per Unit = $ 290

Direct Labour Cost per Unit = $ 340

Variable Manufacturing Overhead per Unit = $ 60

Fixed Manufacturing Overhead per Unit = $ 77.27

Unit Product Cost as per Absorption Costing = 290 + 340 + 60 + 77.27

Unit Product Cost as per Absorption Costing = $ 767.27

Notes

Fixed Manufacturing Overhead per Unit = Fixed Manufacturing Overhead / Total Units Produced

Fixed Manufacturing Overhead = $ 850,000

Total Units Produced = 11,000 Units

Fixed Manufacturing Overhead per Unit = 850,000 / 11,000

Fixed Manufacturing Overhead per Unit = $ 77.27 for One Gamelan

Requirement 2

Calculation of Unit Product Cost as per Variable Costing

Unit Product Cost = Direct Materials Cost per Unit + Direct Labour Cost per Unit + Variable Manufacturing Overhead per Unit

Direct Materials Cost per Unit = $ 290

Direct Labour Cost per Unit = $ 340

Variable Manufacturing Overhead per Unit = $ 60

Unit Product Cost as per Variable Costing = 290 + 340 + 60

Unit Product Cost as per Variable Costing = $ 690 for One Gamelan